# Agent²: An Agent-Generates-Agent Framework for Reinforcement Learning Automation

### v1

@@ Abstract
Reinforcement learning agent development traditionally requires extensive expertise and lengthy iterations, often resulting in high failure rates and limited accessibility. This paper introduces Agent²—a novel agent-generates-agent framework that achieves fully automated RL agent design through intelligent LLM-driven generation. The system autonomously transforms natural language task descriptions and environment code into comprehensive, high-performance reinforcement learning solutions without human intervention.
@@
Agent² features a revolutionary dual-agent architecture. The Generator Agent serves as an autonomous AI designer that analyzes tasks and generates executable RL agents, while the Target Agent is the resulting automatically generated RL agent. The framework decomposes RL development into two distinct stages: MDP modeling and algorithmic optimization, enabling more targeted and effective agent generation. Built on the Model Context Protocol, Agent² provides a unified framework that standardizes intelligent agent creation across diverse environments and algorithms, while incorporating adaptive training management and intelligent feedback analysis for continuous improvement.
@@
Extensive experiments on a wide range of benchmarks—including MuJoCo, MetaDrive, MPE, and SMAC—demonstrate that Agent² consistently outperforms manually designed solutions across all tasks, achieving up to 55% performance improvement and substantial gains on average. By enabling truly end-to-end, closed-loop automation, this work establishes a new paradigm in which intelligent agents design and optimize other agents, marking a fundamental breakthrough for automated AI systems.
Introduction
@@
Reinforcement learning (RL) has achieved remarkable success in diverse domains such as robotics, games, and autonomous systems. However, the process of developing high-performing RL agents remains notoriously complex, requiring substantial domain expertise, intricate environment engineering, careful algorithm selection, and tedious trial-and-error tuning (Li 2017). This manual, expertise-driven workflow has become a critical bottleneck to the practical adoption and democratization of RL, resulting in high entry barriers, limited scalability, and suboptimal research efficiency.
Recent advances in large language models (LLMs) have demonstrated unprecedented capabilities in autonomous reasoning and code synthesis (Achiam et al. 2023). These advancements make it possible to automate not just component-level tasks, but the entire RL agent development pipeline without human intervention.
@@
In this work, we present Agent2, an agent-generates-agent framework that achieves end-to-end automation of RL agent design and optimization using LLMs. Agent2 introduces a novel dual-agent architecture: the Generator Agent acts as an autonomous AI designer capable of analyzing and generating all components required for RL agent; the resulting Target Agent is then constructed from these components and interacts with the environment for training and evaluation.
A key innovation of Agent2 is its intelligent two-stage decomposition of the RL development process. First, Agent2 convert high-level task descriptions into formal MDP representations, capturing environment dynamics, rewards, and constraints. Next, Agent2 automatically selects algorithms, designs neural architectures, and tunes hyperparameters, guided by LLM reasoning and empirical feedback. Integrated adaptive training and feedback analysis further enable robust error correction and iterative performance improvement through closed-loop interaction with the environment.
The remainder of this paper is organized as follows. We review related work, introduce our methodology, describe experimental details and results, and conclude with final remarks.
Related Work
@@
Over the past few years, a wide range of general RL frameworks have been developed, such as RLlib (Liang et al. 2018), Stable Baselines3 (Raffin et al. 2021), Tianshou (Weng et al. 2022), and Xuance (Liu et al. 2023). These frameworks provide standardized implementations of algorithms, unified experiment management, and convenient interfaces for environment integration and distributed training. While these systems have greatly facilitated RL research and application, developing RL agents still heavily depends on expert knowledge and manual intervention. Currently, AutoML (He, Zhao, and Chu 2021) has seen rapid development and become relatively mature, but automated reinforcement learning (AutoRL) remains more complex and is still in its early stages. The ARLO framework (Mussi et al. 2023) systematically modularizes the AutoRL pipeline and supports automated selection and optimization of most stage. However, its automation is limited, as key stages such as feature engineering and data preparation often depend on hand-crafted methods and expert knowledge. Recently, LLMs have enabled new opportunities for RL automation by intelligently generating or optimizing key components of the RL pipeline (Cao et al. 2024; Sun, Huang, and Pompili 2024).
@@
For MDP modeling, some studies have explored automating the design of state representations, reward functions, and action spaces. For state representation, YOLO-MARL (Zhuang et al. 2024) and LERO (Wei et al. 2025) use LLMs to infer high-level or hidden context in multi-agent environment, helping to address limitations from partial observability. In single-agent scenarios, LESR (Wang et al. 2024) and ExplorLLM (Ma et al. 2024a) use LLMs to autonomously generate or reformulate task-related state representations, improving learning efficiency and accelerating training. Reward function design has also been extensively studied in both single-agent and multi-agent RL. In particular, single-agent studies have addressed the challenge of sparse rewards by developing automated reward shaping methods, especially for tasks that require complex robotic control (Ma et al. 2024b; Song et al. 2023) or operate in high-level environments such as Minecraft (Li et al. 2024). In multi-agent settings, research mainly focuses on solving the credit assignment problem for effective reward distribution (Nagpal et al. 2025; Lin et al. 2025; He et al. 2025). For action spaces, recent methods use LLM-generated action masking or suboptimal policies to dynamically constrain and guide RL agents, incorporating user preferences and improving sample efficiency and policy adaptability (Zhao et al. 2025; Karine and Marlin 2025).
@@
Beyond these components, the design of neural network architectures and the optimization of hyperparameters also play a crucial role in RL performance. Recent work has thus explored automating both aspects to further reduce manual effort. LLMatic (Nasir et al. 2024) combines LLM code generation with quality diversity optimization to discover diverse and effective architectures, while EvoPrompting (Chen, Dohan, and So 2023) uses LMs as adaptive operators in evolutionary NAS. In addition to network architecture design, recent studies have also leveraged LLMs to automate hyperparameter optimization in RL. Some studies use LLMs to suggest and iteratively refine hyperparameter configurations based on dataset and model descriptions (Liu, Gao, and Li 2024; Zhang et al. 2023). Additionally, SLLMBO (Mahammadli and Ertekin 2024) introduces a sequential LLM-based framework that adapts the search space dynamically and enhances optimization robustness. These methods highlight the potential of LLMs to design the network architecture and optimize the hyperparameter of RL algorithms.
However, most existing AutoRL approaches automate only a single stage of the RL pipeline, rather than enabling fully end-to-end, closed-loop agent generation. As a result, substantial expert intervention is still required to bridge the gap between task specification and practical agent deployment.
Methods
@@
Figure 1: Agent2 processes natural language task descriptions, environment code, and constraints as input. Through problem analysis, it decomposes the task and assesses its complexity before invoking Model Context Protocol Services, which leverage LLMs for knowledge integration, natural language understanding, performance analysis, and code synthesis. The workflow then proceeds to MDP modeling—including state space, action space, and reward function design—followed by algorithmic optimization, where the agent autonomously selects algorithms, designs network architectures, and tunes hyperparameters. The generated components are assembled into the Target Agent, which is ready for training and evaluation. The entire process supports iterative refinement to enhance solution quality.
Task-to-MDP Mapping
Automatic Problem Analysis
A key prerequisite for automated RL agent generation is the precise understanding and formalization of the target problem. To address this, Agent2 begins the RL pipeline with a dedicated problem analysis stage, systematically bridging the gap between high-level user intent and formal RL formulation.
The Generator Agent leverages LLMs to process heterogeneous user input, typically including a natural language task description and, when available, environment code or specifications. The LLM integrates these inputs with its internal and, if configured, external knowledge bases, producing a structured representation of the RL problem. This analysis enables the system to: Identify learning objectives and reward structure from the task description; Infer state, action, and transition dynamics from the environment code; Extract explicit and implicit constraints, such as safety or operational limits; Anticipate challenges such as partial observability or delayed rewards.
@@
MDP Modeling
The MDP modeling module automates the refinement of initial, suboptimal RL environments into effective MDPs for reinforcement learning. In practice, available simulators or environments may only provide a coarse or incomplete MDP specification, requiring expert effort to redesign the observation space, action space, or reward function. This module leverages LLMs to analyze and optimize these core components, minimizing manual intervention. While LLMs are capable of constructing MDPs from scratch, our approach focuses on the more practical and challenging scenario of automatically optimizing an existing but imperfect environment.
@@
The LLM examines environment code and task objectives to identify and refine relevant state variables, possibly applying feature selection, dimensionality reduction, or combining observations to improve learning efficiency. The refined observation space is expressed as $s^{\prime}=f_{\text{obs}}(s)\in\mathcal{S}^{\prime}$, where $f_{\text{obs}}$ denotes the learned or synthesized transformation function.
Similarly, the action space is optimized to balance expressiveness and simplicity, with the LLM recommending discretization, merging, splitting, or reparameterization as needed: $a^{\prime}=f_{\text{act}}(a)\in\mathcal{A}^{\prime}$, with $f_{\text{act}}$ constructed through automatic analysis and synthesis.
@@
Each component is encapsulated as a wrapper function, enabling seamless integration and algorithmic verification within the RL pipeline.
Adaptive Verification and Refinement
@@
This module aims to ensure that all synthesized components are executable, and optimized for agent learning. Each component—observation, action, and reward functions ($f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$)—is first integrated through standardized interfaces and checked by a verification operator $V$. If verification fails, an error feedback $e$ is generated and, together with the original RL analysis $L_{\text{analysis}}$ and component $f$, forms an adaptive prompt $P_{\text{error}}$. The LLM inference function $F_{\text{LLM}}$ then synthesizes a revised component $f^{*}$:
$\text{If}\;V(f)=\text{False},\;\text{then}\;f^{*}\leftarrow F_{\text{LLM}}\big{(}P_{\text{error}}(e,L_{\text{analysis}},f)\big{)}$ (2)
This process repeats until all components pass verification. Beyond static verification, the system continuously monitors the agent’s training dynamics and summarizes key performance indicators from TensorBoard data into a concise report, which serves as the performance feedback $\epsilon$ for further optimization. This feedback, together with $L_{\text{analysis}}$, $f$, and training history $\mathcal{H}$, to construct a prompt $P_{\text{perf}}$. The LLM then produces improved components:
$f^{*}\leftarrow F_{\text{LLM}}\big{(}P_{\text{perf}}(\epsilon,L_{\text{analysis}},f,\mathcal{H})\big{)}$ (3)
This closed-loop process ensures that MDP components are robust, reliable, and continually optimized, advancing the automation and practicality of LLM-based RL agent development.
@@
Algorithm 1 Task to MDP Mapping Pipeline
0: Task description $T_{\text{task}}$, environment information $T_{\text{env}}$, additional constraints $T_{c}$, LLM inference function $F_{\text{LLM}}$
0: Verified and optimized MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$
1:// Automatic Problem Analysis
[…]
3:// MDP Modeling
[…]
5:// Adaptive Verification and Refinement
6: Initialize best components $f^{*}\leftarrow f^{(0)}$, best score $S^{*}_{\text{hist}}\leftarrow S^{(0)}$, training history $\mathcal{H}\leftarrow\{\}$
7:for $t=1$ to $T$ do
8:  Integrate $f^{(t)}$ into RL pipeline
9:  Apply verification operator $V$ to $f^{(t)}$
10:  if verification fails (error eethen
11:   $f^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{error}}(e,L_{\text{analysis}},f^{(t)}))$
12:   continue
13:  end if
14:  Train agent with $f^{(t)}$, collect TensorBoard metrics, update $\mathcal{H}$
15:  Summarize key metrics into feedback $\epsilon^{(t)}$
16:  Obtain performance score $S^{(t)}$
17:  if $S^{(t)}>S^{*}_{\text{hist}}$ then
18:   $S^{*}_{\text{hist}}\leftarrow S^{(t)}$, $f^{*}\leftarrow f^{(t)}$
19:  end if
20:  if not converged then
21:   $f^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{refine}}(\epsilon^{(t)},L_{\text{analysis}},f^{(t)},\mathcal{H}))$
22:  else
23:   break
24:  end if
25:end for
26:return  Final verified and optimized MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$
Algorithmic Optimization
@@
Agent2 uses LLMs to adaptively select algorithms based on comprehensive information, including the structured RL analysis $L_{\text{analysis}}$, environment code $T_{\text{env}}$, and current MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$. Given a set of candidates $G$, the LLM identifies the most suitable algorithm $g^{*}\in G$ and provides a brief rationale $L_{\text{algo}}$. Formally, the prompt is $P_{\text{alg}}(L_{\text{analysis}},f,T_{\text{env}})$, and the LLM inference function performs:
$g^{*},\ L_{\text{algo}}=F_{\text{LLM}}\big{(}P_{\text{alg}}(G,f,L_{\text{analysis}},T_{\text{env}})\big{)}$ (4)
Network Architecture Design
@@
The Generator Agent consolidates structured information from the MDP model ($f_{\text{obs}}$, $f_{\text{act}}$, $f_{\text{rew}}$), network $N^{*}$, selected algorithm $g^{*}$, reference configurations $J_{\text{hp}}$, and training history $\mathcal{H}$. Prompt engineering guides the LLM to assess current settings, identify bottlenecks, and propose incremental, interpretable adjustments prioritized for training stability and performance.
@@
$Hp^{*},\ L_{\text{hp}}=F_{\text{LLM}}\Big{(}P_{\text{hp}}(g^{*},N^{*},f_{\text{obs}},f_{\text{act}},f_{\text{rew}},J_{\text{hp}},\mathcal{H})\Big{)}$ (6)
@@
Configuration Integration and Refinement

### v2

@@ Abstract
Reinforcement learning (RL) agent development traditionally requires substantial expertise and iterative effort, often leading to high failure rates and limited accessibility. This paper introduces Agent2, an LLM-driven agent-generates-agent framework for fully automated RL agent design. Agent2 autonomously translates natural language task descriptions and environment code into executable RL solutions without human intervention.
@@
The framework adopts a dual-agent architecture: a Generator Agent that analyzes tasks and designs agents, and a Target Agent that is automatically generated and executed. To better support automation, RL development is decomposed into two stages—MDP modeling and algorithmic optimization—facilitating targeted and effective agent generation. Built on the Model Context Protocol, Agent2 provides a unified framework for standardized agent creation across diverse environments and algorithms, incorporating adaptive training management and intelligent feedback analysis for continuous refinement.
@@
Extensive experiments on benchmarks including MuJoCo, MetaDrive, MPE, and SMAC show that Agent2 outperforms manually designed baselines across all tasks, achieving up to 55% performance improvement with consistent average gains. By enabling a closed-loop, end-to-end automation pipeline, this work advances a new paradigm in which agents can design and optimize other agents, underscoring the potential of agent-generates-agent systems for automated AI development.
1 Introduction
@@
Reinforcement learning (RL) has achieved remarkable success in diverse domains such as robotics, games, and autonomous systems. However, the process of developing high-performing RL agents remains notoriously complex, requiring substantial domain expertise, intricate environment engineering, careful algorithm selection, and tedious trial-and-error tuning (Li, 2017). This manual, expertise-driven workflow has become a major barrier to the practical adoption and broader accessibility of RL, leading to high entry costs, limited scalability, and reduced research efficiency. Recent advances in large language models (LLMs) have demonstrated unprecedented capabilities in autonomous reasoning and code synthesis (Achiam et al., 2023). These advancements make it possible to automate not just component-level tasks, but the entire RL agent development pipeline without human intervention.
@@
In this work, we present Agent2, an agent-generates-agent framework that introduces a novel dual-agent architecture. The Generator Agent serves as an autonomous AI designer, capable of analyzing and producing all necessary components for an RL agent. Based on these components, the Target Agent is constructed and subsequently interacts with the environment for training and evaluation.
A key innovation of Agent2 lies in its end-to-end automated pipeline for RL agent generation. Instead of merely automating isolated components, Agent2 achieves unified and functionalized development of each module, efficient information exchange across modules, and integrated training frameworks. Moreover, Agent2 incorporates mechanisms for detecting and resolving training anomalies, enabling iterative evolution and continual refinement. This comprehensive design transforms RL development into a adaptive and progressively improving pipeline. The remainder of this paper is organized as follows. Section 2 reviews related work, Section 3 presents our methodology, Section 4 reports experimental results, and Section 5 concludes the paper.
2 Related Work
@@
Over the past few years, a wide range of general RL frameworks have been developed, such as RLlib (Liang et al., 2018), Tianshou (Weng et al., 2022), and Xuance (Liu et al., 2023). These frameworks provide standardized implementations of algorithms, unified experiment management, and convenient interfaces for environment integration and distributed training. While these frameworks have greatly facilitated RL research and application, developing RL agents still heavily depends on expert knowledge and manual intervention. LLMs have enabled new opportunities for RL automation by intelligently generating or optimizing key components of the RL pipeline (Cao et al., 2024; Sun et al., 2024).
@@
We contend that the RL automation can be broadly categorized into two complementary dimensions: MDP modeling and algorithmic optimization. For MDP modeling, some studies have explored automating the design of state representations, reward functions, and action spaces. LESR (Wang et al., 2024) and ExplorLLM (Ma et al., 2024a) use LLMs to autonomously generate or reformulate task-related state representations, improving learning efficiency and accelerating training. YOLO-MARL (Zhuang et al., 2024) and LERO (Wei et al., 2025) use LLMs to infer high-level or hidden context in multi-agent environment, helping to address limitations from partial observability. Some studies of reward function design have addressed the challenge of sparse rewards by developing automated reward shaping methods, especially for tasks that require complex robotic control (Ma et al., 2024b; Song et al., 2023) or operate in high-level environments such as Minecraft (Li et al., 2024). In multi-agent settings, research mainly focuses on solving the credit assignment problem for effective reward distribution (Nagpal et al., 2025; Lin et al., 2025; He et al., 2025). For action spaces, recent methods use LLM-generated action masking or suboptimal policies to dynamically constrain and guide RL agents, incorporating user preferences and improving sample efficiency and policy adaptability (Zhao et al., 2025; Karine & Marlin, 2025).
@@
Research on its algorithmic optimization mainly follows the AutoML, which has seen rapid development and become relatively mature (He et al., 2021). Algorithmic optimization focusing on two directions: neural network architecture design and hyperparameter optimization. LLMatic (Nasir et al., 2024) combines LLM code generation with quality diversity optimization to discover diverse and effective architectures, while EvoPrompting (Chen et al., 2023) uses LMs as adaptive operators in evolutionary NAS. For hyperparameter optimization, some studies use LLMs to suggest and iteratively refine hyperparameter configurations based on dataset and model descriptions (Liu et al., 2024; Zhang et al., 2023; Mussi et al., 2023). Additionally, SLLMBO (Mahammadli & Ertekin, 2024) introduces a sequential LLM-based framework that adapts the search space dynamically and enhances optimization robustness. These methods highlight the potential of LLMs to design the network architecture and optimize the hyperparameter of RL algorithms.
However, most existing AutoRL approaches automate only a single stage of the RL pipeline, rather than enabling fully end-to-end agent generation, which requires the coordination of multiple modules as well as the handling of their interactions and debugging. This poses a far more challenging problem, which we address in the following section.
3 Methods
@@
Figure 1: The framework of Agent2 consists of three main stages. Firstly, Agent2 analyzes the problem using natural language task descriptions and environment code as inputs. Secondly, the framework proceeds to MDP modeling, including the design of the state space, action space, and reward function. Thirdly, the framework is followed by algorithmic optimization, where the agent autonomously selects appropriate algorithms, designs network architectures, and tunes hyperparameters. The entire framework operates in compliance with the Model Context Protocol (MCP), ensuring standardized integration of services. Finally, the generated components are assembled into the Target Agent, which is ready for training and evaluation. The entire process supports iterative refinement to enhance solution quality.
3.1 Task-to-MDP Mapping
3.1.1 Automatic Problem Analysis
A key prerequisite for automated RL agent generation is the precise understanding and formalization of the target problem. The Generator Agent leverages LLMs to process heterogeneous user input with its knowledge bases, producing a structured representation of the RL problem. This analysis enables the system to: Identify learning objectives and reward structure from the task description; Infer state, action, and transition dynamics from the environment code; Extract explicit and implicit constraints, such as safety or operational limits; Anticipate challenges such as partial observability or sparse rewards.
@@ 3.1.1 Automatic Problem Analysis
3.1.2 MDP Modeling
The MDP modeling module automates the refinement of initial RL environments into effective MDPs for reinforcement learning. In practice, available simulators or environments may only provide a coarse or incomplete MDP specification, requiring expert effort to redesign the observation space, action space, or reward function. This module leverages LLMs to analyze and optimize these core components, minimizing manual intervention. While LLMs are capable of constructing MDPs from scratch, our approach focuses on the more practical and challenging scenario of automatically optimizing an existing but imperfect environment.
@@ 3.1.2 MDP Modeling
The LLM examines environment code and Problem Analysis to identify and refine relevant state variables, possibly applying feature selection, dimensionality reduction, or combining observations to improve learning efficiency. The refined observation space is expressed as $s^{\prime}=f_{\text{obs}}(s)\in\mathcal{S}^{\prime}$, where $f_{\text{obs}}$ denotes the learned or synthesized transformation function.
Similarly, the action space is optimized to balance expressiveness and simplicity, with the LLM recommending merging, splitting, or reparameterization as needed: $a^{\prime}=f_{\text{act}}(a)\in\mathcal{A}^{\prime}$, with $f_{\text{act}}$ constructed through automatic analysis and synthesis.
@@
Each component is encapsulated as a wrapper function, enabling seamless integration and algorithmic verification within the RL framework.
3.1.3 Adaptive Verification and Refinement
@@
Each component—observation, action, and reward functions ($f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$)—is first integrated through standardized interfaces and checked by a verification operator $V$. If verification fails, an error feedback $e$ is generated and, together with the original RL analysis $L_{\text{analysis}}$ and component $f$, forms an adaptive prompt $P_{\text{error}}$. The LLM inference function $F_{\text{LLM}}$ then synthesizes a revised component $f^{*}$.
Beyond static verification, the system continuously monitors the agent’s training dynamics and summarizes key performance indicators from TensorBoard data into a concise report, which serves as the performance feedback $\epsilon$ for further optimization. This feedback, together with $L_{\text{analysis}}$, $f$, and training history $\mathcal{H}$, to construct a prompt $P_{\text{perf}}$. The LLM then produces improved components. This closed-loop process ensures that MDP components are robust, reliable, and continually optimized, advancing the automation and practicality of LLM-based RL agent development.
@@
Algorithm 1 Adaptive Verification and Refinement of MDP Components
0: Initial MDP components $f^{(0)}=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$, LLM inference function $F_{\text{LLM}}$, RL analysis $L_{\text{analysis}}$, maximum iterations $T$
0: Verified and optimized MDP components $f^{*}$
1: Initialize best components $f^{*}\leftarrow f^{(0)}$, best score $S^{*}_{\text{hist}}\leftarrow S^{(0)}$, training history $\mathcal{H}\leftarrow\{\}$
2:for $t=1$ to $T$ do
3:  Integrate $f^{(t)}$ into RL pipeline
4:  Apply verification operator $V$ to $f^{(t)}$
5:  if $V(f^{(t)})=\text{False}$ with error $e$ then
6:   $f^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{error}}(e,L_{\text{analysis}},f^{(t)}))$
7:   continue
8:  end if
9:  Train agent with $f^{(t)}$, collect TensorBoard metrics $\epsilon^{(t)}$, obtain score $S^{(t)}$, update $\mathcal{H}$
10:  if $S^{(t)}>S^{*}_{\text{hist}}$ then
11:   $S^{*}_{\text{hist}}\leftarrow S^{(t)}$, $f^{*}\leftarrow f^{(t)}$
12:  end if
13:  if not converged then
14:   $f^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{perf}}(\epsilon^{(t)},L_{\text{analysis}},f^{(t)},\mathcal{H}))$
15:  else
16:   break
17:  end if
18:end for
19:return $f^{*}$
3.2 Algorithmic Optimization
@@ 3.2.1 Algorithm Selection
Agent2 uses LLMs to adaptively select algorithms based on comprehensive information, including the structured RL analysis $L_{\text{analysis}}$, environment code $T_{\text{env}}$, current MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$, and training history $\mathcal{H}$. Given a set of candidates $G$, the LLM identifies the most suitable algorithm $g^{*}\in G$ and provides a brief rationale $L_{\text{algo}}$. Formally, the LLM inference function performs:
$g^{*},\ L_{\text{algo}}=F_{\text{LLM}}\big(P_{\text{alg}}(G,f,L_{\text{analysis}},T_{\text{env}},\mathcal{H})\big)$ (2)
3.2.2 Network Architecture Design
@@ 3.2.3 Hyperparameter Optimization
The Generator Agent consolidates structured information from the MDP model $f^{*}$, network $N^{*}$, selected algorithm $g^{*}$, reference configurations $J_{\text{hp}}$, and training history $\mathcal{H}$. Prompt engineering guides the LLM to assess current settings, identify bottlenecks, and propose incremental, interpretable adjustments prioritized for training stability and performance.
@@
$Hp^{*},\ L_{\text{hp}}=F_{\text{LLM}}\Big(P_{\text{hp}}(g^{*},N^{*},f^{*},J_{\text{hp}},\mathcal{H})\Big)$ (4)
@@
3.2.4 Configuration Integration and Refinement